OpenStudy (anonymous):

Explain the meaning of a 2 by 3 matrix. Give an example from daily life of data you could organize in a matrix.

OpenStudy (anonymous):

2X3 matrix looks like this: \[\left[\begin{matrix}a & b & c \\ d& e&f\end{matrix}\right]\] It usually represents two equations in three variables: \[\left[\begin{matrix}a & b & c \\ d& e&f\end{matrix}\right]\left(\begin{matrix}x \\ y \\ z\end{matrix}\right)=\left(\begin{matrix}l \\ m \end{matrix}\right)\] Means: \[ax+by+cz=l\] \[dx+ey+fz=m\]